An object traveling with a speed of 10 π‘š/𝑠 slides on a horizontal floor. How far will it travel before coming to rest if the coefficient of friction is 0.30?


1
Expert's answer
2021-10-01T12:22:37-0400

"F=ma\to \mu mg=ma\to a=\mu g=0.3\cdot9.81=2.94\ (m/s^2)"


"s=\frac{v_0^2}{2a}=\frac{10^2}{2\cdot 2.94}=17\ (m)" . Answer